OpenStudy (anonymous):

What is the square root of minus 1. When i do it in a scientific calculator, it says 'i' (imaginary unit) and i have no idea what that means

OpenStudy (anonymous):

\[\sqrt{-1} = x \] what is the value of x

OpenStudy (anonymous):

please dont put links

OpenStudy (anonymous):

Imaginary numbers result when you are taking the square root of any negative number: For an example, \[\sqrt{-1}\] = i \[\sqrt{-25} = 5i\] The 'i' is kind of like the negative sign

OpenStudy (anonymous):

so what is an imaginary unit

OpenStudy (anonymous):

I know of a video you can watch if you want to

OpenStudy (anonymous):

yes what is the vid @zaynahf

OpenStudy (hba):

did you see the link @awn786 It has everything.

OpenStudy (anonymous):

yes i did but what is an imaganary unit

OpenStudy (anonymous):

WOW OMG complicated @hba

OpenStudy (anonymous):

The first link is really straight forward, try that

OpenStudy (hba):

@awn786 Well,I tried to simplify and put everything :)

OpenStudy (anonymous):

thanks for trying. @hba

OpenStudy (hba):

You are welcome :)

OpenStudy (anonymous):

listen and can u give a simple i word answer: \[\sqrt{-1}=x\] what is the value of x in numbers.

OpenStudy (anonymous):

1 word answer*

OpenStudy (anonymous):

not i word

OpenStudy (hba):

I can help you understand what imaginary numbers are. So,the base of imaginary numbers is actually laid upon some of the things like \[\huge \ i=\sqrt{-1}\] The other thing is \[\huge\ \omega=\frac{ -1+i \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

whats w

OpenStudy (hba):

We call this omega,It's derivations is given in my link :)

OpenStudy (anonymous):

whats w @hba

OpenStudy (hba):

Now as per your question. x=root{-1} x=i Here i is actually iota and it is an imaginary number.

OpenStudy (hba):

That's omega @awn786

OpenStudy (anonymous):

whats omega

OpenStudy (hba):

@awn786 Please check the link i have provided. It has everything Please give it a read.

OpenStudy (anonymous):

i said in numbers. your answer is basicly my question

OpenStudy (hba):

Well,In numbers it is sqrt{-1}

OpenStudy (anonymous):

it should be easy shouldent it.

OpenStudy (anonymous):

i dont understand ur link

OpenStudy (anonymous):

what is omega

OpenStudy (hba):

I have already told you what omega is.

OpenStudy (anonymous):

no u havent @hba

OpenStudy (hba):

I said it is -1+i sqrt{3}/2

OpenStudy (anonymous):

ok (i dont get it but ok) i' only in year 9 doing my gcse's

OpenStudy (anonymous):

\[\sqrt{-1}\] doesn't exist as a the REAL number you are asking for, because there's no number that you can square (because of the SQUARE root) and gives you -1

OpenStudy (anonymous):

you just need to know what a square root is

OpenStudy (anonymous):

i know what a square root is ofc

OpenStudy (anonymous):

that's why there's no sqrt of ANY negative numbers, because everthing squared is BIGGER than 0

OpenStudy (anonymous):

(possitive)

OpenStudy (anonymous):

oh i get it so its an imaganary number because it cant exist?

OpenStudy (anonymous):

you have the best responce

OpenStudy (anonymous):

EXACTLY, it CAN'T EXIST as a "normal number"

OpenStudy (anonymous):

thank u so much.

OpenStudy (anonymous):

^^^^^^ @GCR92

OpenStudy (anonymous):

so mathematicians around XVII century had to figure out another type of numbers, called COMPLEX numbers, and the first number of complex numbers is \[\sqrt{-1}\]

OpenStudy (anonymous):

they INVENTED it

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

ur a genius @GCR92

OpenStudy (anonymous):

not really, but thanks!

OpenStudy (anonymous):

ill draw you something

OpenStudy (anonymous):

you can bring the most complex thing and make others understand. are u sure ur not related to prfsr. brian cox

OpenStudy (anonymous):